OpenStudy (anonymous):

A 75.0 kg astronaut is training for accelerations that he will experience upon reentry. He is placed in a centrifuge (r = 8.0 m) and spun at a constant angular velocity of 25.3 rpm. Answer the following: e.What is the centripetal force and net torque experienced by the astronaut? Give magnitudes and directions.

OpenStudy (anonymous):

a = (v^2)/r T = m(r^2)a